Find and the solutions of the following inequalities.
Express your answer using interval notation.
1.5x-12 (is less than or equal to) 2x

Answer by checkley77(12844)

1.5x-12<=2x
1.5x-2x<=12
-.5x<=12
x>=12/-.5 (note the sign change when dividing by a negative value)
x>=-24 ans.
Proof:
Let x=-23
1.5*-23-12<2*-23
-34.5-12<=-46
-46.5<=-46